### v1

# QuantumBind-RBFE: Accurate Relative Binding Free Energy Calculations Using Neural Network Potentials

Abstract
Accurate prediction of protein-ligand binding affinities is crucial in drug discovery, particularly during hit-to-lead and lead optimization phases, however, limitations in ligand force fields continue to impact prediction accuracy. In this work, we validate relative binding free energy (RBFE) accuracy using neural network potentials (NNPs) for the ligands. We utilize a novel NNP model, AceForce 1.0, based on the TensorNet architecture for small molecules that broadens the applicability to diverse drug-like compounds, including all important chemical elements and supporting charged molecules. Using established benchmarks, we show overall improved accuracy and correlation in binding affinity predictions compared with GAFF2 for molecular mechanics and ANI2-x for NNPs. Slightly less accuracy but comparable correlations with OPLS4. We also show that we can run the NNP simulations at 2 fs timestep, at least two times larger than previous NNP models, providing significant speed gains. The results show promise for further evolutions of free energy calculations using NNPs while demonstrating its practical use already with the current generation. The code and NNP model are publicly available for research use.
[…]
1 Introduction
Accurate prediction of protein-ligand binding affinities is crucial in drug discovery, particularly in the hit-to-lead and lead optimization stages, where a congeneric series of ligands must be screened efficiently. Among the methods employed, alchemical free energy calculations1, 2, 3, 4, 5, including relative binding free energy (RBFE) methods, have gained prominence due to their capacity to deliver reliable estimates of binding affinities across a wide range of compounds. Such techniques are widely used in both academia and the pharmaceutical industry. However, the accuracy of RBFE calculations is hampered by several factors such as convergence of the simulations, omission of relevant chemical effects (e.g. tautomerizations and/or protonation), poor selection of pair edges, poor pose selection and/or the inaccuracy of the ligand force field. In this work, we put a special focus on this last point.
Accurately capturing molecular interactions across the diverse landscape of drug-like compounds is a challenge for traditional molecular mechanics (MM) force fields like GAFF6, 7, CGenFF8, 9, and OpenFF10, 11. These force fields can struggle with rare chemical groups and cannot account for key energetic factors, such as polarization and quantum effects. As a result, they can introduce inaccuracies in binding free energy predictions, especially when handling complex ligand chemistry or significant conformational flexibility.
Recently, an alternative approach has emerged based on neural network potentials (NNPs) to use fast neural network function approximation of the quantum mechanic’s energy surface12. Based on these architectures, a few methods have been made available which we will consider in this work. MACE-based models13 are very accurate but computationally expensive. This is not a problem in material science where the alternative is very expensive quantum chemistry methods. For biology-oriented applications, ANI-models14 are much faster but at the expenses of accuracy, still providing a very good trade-off. AIMNet215 extends ANI with a different architecture to more elements and molecules, still adding on computational costs. The equivariant transform (ET) architecture16 and TensorNet17 aimed to provide state-of-the-art accuracy without sacrificing computational efficiency, being taylored specifically for drug discovery. These two are integrated in the TorchMD-Net software framework 18, 16, 19 and a new NNP model has been made available, AceForce20. Molecular dynamics (MD) simulations based on NNPs promise accurate simulations at reasonable costs compared with density functional theory (DFT) methods, but do represent an increase in computational costs compared to traditional molecular mechanics, yet their future is very promising 21.
Recently, we explored the integration of NPs into RBFE calculations with promising results.22 Specifically, we used ANI-2x14 with good results. However, ANI-2x cannot handle charged species and has a restricted range of supported atom elements. These limitations restrict the applicability across the diverse chemical space encountered in drug discovery. In this work, we extend our previous work22 by testing QuantumBind-RBFE, our NNP/MM approach23 for RBFE calculations, using AceForce. The AceForce 1.020 model supports a broad range of atom elements, including charged molecules, and is tailored for accurate and efficient RBFE calculations across diverse molecular systems. To validate the performance of RBFE calculations using NNPs, we conducted a comprehensive benchmarking study using established datasets, including both charged and neutral ligands. Our results show that we can improve the accuracy of RBFE calculations compared to traditional force fields and our previous test with ANI-2x, reaching state-of-the-art correlations.
2 Methods
2.1 RBFE calculation within the NNP/MM scheme
Calculations were performed using an NNP/MM approach, which combines neural network potentials (NNP) for high-accuracy modeling of ligand interactions and molecular mechanics (MM) for the remainder of the system.23 This hybrid method allows the ligand to be simulated with the more accurate neural network potential, while the surrounding protein environment is treated with classical molecular mechanics for speed.(Figure 1) This setup is particularly advantageous for capturing the finer details of intra-molecular ligand interactions that are typically challenging for traditional MM force fields, like ligand strain.
In NNP/MM, the system is divided into NNP and MM regions, akin to the partitioning used in QM/MM simulations. The total potential energy ($V$) is calculated as:
$V(\vec{r})=V_{\text{NNP}}(\vec{r}_{\text{NNP}})+V_{\text{MM}}(\vec{r}_{\text{MM}})+V_{\text{NNP-MM}}(\vec{r}),$ (1)
where $V_{\text{NNP}}$ and $V_{\text{MM}}$ represent the potential energies for the NNP and MM regions, respectively, and $V_{\text{NNP-MM}}$ accounts for the coupling between these regions. The NNP potential is computed based on the atomic positions and charges within the NNP region, ensuring smooth interaction with the MM-treated environment. 23
Figure 1: Description of the NNP/MM scheme. While the ligand is simulated with a neural network potential (NNP), the rest of the system is treated with classical molecular mechanics (MM)
We utilized the Alchemical Transfer Method (ATM) to perform the RBFE calculations.24 This methodology has been previously validated across multiple benchmarks,25, 26 demonstrating its reliability for accurate binding free energy predictions. Notably, ATM was also employed in our recent work validating its use with NNPs, specifically ANI2x.22 In this study, we extend the previous work substantially.
2.2 Dataset selection
We selected a widely recognized public benchmark dataset that serves as a reference for RBFE studies. The BACE, CDK2, JNK1, MCL1, P38, THROMBIN, and TYK2 targets are part of Wang et al.’s dataset, commonly known as the ”JACS dataset” or ”Schrödinger dataset” 27. We had to omit PTP1B from the evaluation study since all ligands in the series have charge -2 and AceForce, for now, can only compute ligands with charges -1,0 and +1. Compared to our previous work, we can now perform the benchmark in completion. This dataset provides a total of 7 protein targets, 179 ligands, and 280 edges. Comparisons will be made with GAFF2 and FEP+ with the OPLS4 force field28.
2.3 NNP model
AceForce 1.0 is the first version of a new family of potentials 20. It uses TensorNet 1-layer 17 trained on Acellera29’s internal proprietary dataset of forces and energies computed the wB97M-V/def2-tzvppd level of theory. This work is the first test of this model for accuracy in RBFE calculations. A basic evaluation of the model on the Seller’s torsion scan 30 benchmark shows promising performance (Figure 2). For all molecules in the test, we performed the full torsion scans, minimizing the geometry with constrained torsion angles using the forces and energy from each calculation method. For comparison, we did the same procedure with ANI-2x14 from TorchANI31, xTB32, AIMNet215, and MACE-OFF13. We used GeomeTRIC33 for the constrained optimizations. The MAE between the curves for each molecule is shown in Figure 2. The AceForce model is among the best performing and comparable to MACE but at a much faster computational speed. Also, AceForce is using a faster TensorNet 1-layer model, instead to the 2-layer model generally used in 17 for maximum accuracy.
Figure 2: Seller’s torsion scan benchmark, comparing torsion accuracy against a high level quantum chemistry approximation.
2.4 QuantumBind-RBFE Calculation details
The workflow in this project follows the same procedure as in our previous work22. Ligands were parameterized with GAFF26, 7 with RESP charges, which are calculated with the OpenFF-Recharge package.34 The topologies were generated using the parameterize35 tool. Next, complex systems were prepared with HTMD,36 and atom indexes for ligand alignment were performed, selecting three reference atoms for each ligand. For more information on the selection of these atoms please check previous publications.24, 25.
The energy minimization, thermalization, and equilibration steps followed the procedures described in our previous work.25. RBFE simulations were run in triplicate for each edge for an ensemble of 70 ns per replica, a similar amount of simulation time to our previous22 and similar37 studies. We used the Amber ff14SB parameters38, 39 as well as the TIP3P water model. Classical RBFE simulations were run at a $4$ fs timestep while the NNP/MM runs were run at $2$ fs timestep. Previously, we already evaluated how the accuracy difference between 1fs and 4fs runs is minimal for MM runs.22 In the results section we discuss the reasoning behind selecting such a timestep for the NNP/MM runs. All calculations were run on GPUGRID.net40. The parallel replica exchange molecular dynamics simulations were conducted using QuantumBind-RBFE our implementation of ATM, HTMD36 and the ATMForce potential from OpenMM 8.1.141.
2.5 Analysis
Binding free energies and their corresponding uncertainties from the perturbation energy samples were computed using the Unbinned Weighted Histogram Analysis Method (UWHAM).42 The resulting relative binding free energies ($\Delta\Delta G$) were averaged over the three repeats and used as the predicted value, and the standard deviation across the repeats was used as an error estimate. Absolute $\Delta G$ values were computed via a maximum likelihood estimator with cinnabar for the obtained $\Delta\Delta G$ values.43 The accuracy of the methods was evaluated with several metrics, both from an error and ranking perspective. Relative and absolute binding free energies were compared to experimental measurements of mean absolute error (MAE), root mean square error (RMSE), and Kendall Tau correlation coefficient. Additionally, we evaluated the methods’ ability to prioritize the most active compounds in each dataset by analyzing top compound identification. Specifically, we examined whether the methods correctly identified the top 5 ligands and the top 30% of ligands ranked by experimental binding affinity. For the top 5 metrics, we assessed how many of the five most active compounds were included in the predicted top 5. For the top 30% metric, we determined the percentage of ligands within the experimental top 30% that were also present in the predicted top 30%. All results are reported with 95% confidence intervals, calculated using 1000 bootstrap samples. To ensure a fair comparison and minimize discrepancies that can appear from differences in algorithm implementation, we recalculated the metrics and converted $\Delta G$ values for OPLS4 with FEP+ using our analysis scripts. These calculations were performed by incorporating the $\Delta\Delta G$ values prior to the cycle closure correction into the algorithm. The initial structures and the results of all RBFE simulations are available on GitHub (refer to the Data Availability section for further details).
3 Results
3.1 QuantumBind-RBFE calculations
Table S1 presents the RMSE, MAE, and Kendall tau correlation values for the evaluated targets, comparing GAFF2, OPLS4, and AceForce 1.0. These metrics provide insight into both the accuracy (via RMSE and MAE) and ranking performance (via Kendall tau) of each method across a diverse set of protein-ligand systems. $\Delta\Delta G$ values for the aforementioned systems and methods are included in the supporting information (Supporting TableS7 and Supporting FigureS1. Since this work uses a different charge model (RESP) compared to our previous publications (AM1BCC), we compared the results obtained here with those from earlier studies (Supporting Tables S3, S5, S4). For most targets, the results are very similar between the two charge models. Notably, in certain cases, the improvement with RESP charges is considerable. While the limitations and possible improvements of this approach are discussed in later sections, the overall improvement observed with RESP charges motivated their use for the QuantumBind-RBFE calculations presented in this work.
QuantumBind-RBFE demonstrates improved performance compared to GAFF2 across most targets. When considering all computed data points, AceForce 1.0 achieves better overall metrics, with a notable decrease in RMSE from 1.17 kcal/mol for GAFF2 to 0.99 kcal/mol for AceForce, and a reduction in MAE from 0.90 to 0.79 kcal/mol. The Kendall tau correlation is slightly better for AceForce 1.0 (0.59 vs. 0.55 for GAFF2), though the difference is relatively modest. Importantly, AceForce consistently achieves a mean absolute error (MAE) below 1.0 kcal/mol for all targets, underscoring its robust accuracy across diverse systems.
Despite the similar accuracy in terms of RMSE and MAE across many targets, AceForce 1.0 consistently outperforms GAFF2 in ranking capabilities. Kendall tau correlations are higher for all targets except MCL1 and THROMBIN. The lower correlation for THROMBIN can be attributed to the small size of the dataset (N = 11), where even a single outlier can disproportionately affect the correlation metric, as observed in previous studies.44 For MCL1, the reduced Kendall tau (0.28 for AceForce vs. 0.44 for GAFF2) is primarily due to the overprediction of binding affinities for five ligands in the series, which significantly impacts ranking performance.
AceForce 1.0 demonstrates competitive performance compared to the reported results with OPLS4,28 though the latter generally achieves lower RMSE and MAE values across most targets. When considering all computed data points, OPLS4 outperforms AceForce 1.0 with a lower RMSE (0.78 vs. 0.99 kcal/mol) and MAE (0.61 vs. 0.79 kcal/mol). Kendall tau correlations are also slightly higher for OPLS4 (0.66 vs. 0.59 for AceForce), indicating better overall ranking accuracy.
Overall, OPLS4 show still superior RMSE for most targets, while comparable ranking performance to AceForce 1.0 where some targets are better ranked by OPLS4, some by AceForce 1.0. It is worth noting that the methodologies for RBFE calculations differ between QuantumBind and FEP+, including distinct protocols and molecular dynamics engines. These differences can introduce additional variability in the results. The observed differences are also usually small and often within the margin of error, emphasizing the promise of NNPs as a viable alternative for RBFE calculations.
Finally, we highlight the significant improvement AceForce 1.0 provides over GAFF2 when using the same protocol and molecular dynamics engine. This underscores the robustness of AceForce 1.0 as an advanced tool for RBFE predictions in drug discovery workflows. The parameterization of small molecules with NNPs, even using the NNP/MM scheme, becomes trivial, as it does not require the calculation and fitting of torsion scans, greatly simplifying the process.
Figure 3: (Top) Root Mean Squared Error (RMSE) and (bottom) Kendall tau correlation for the $\Delta G$s of each protein-ligand system calculated in combination with different approaches: GAFF2 (teal), reported estimates using FEP+ with the OPLS4 forcefield(yellow) and AceForce 1.0 (blue).
Figure 4: Scatterplots of predicted $\Delta G$ values for each evaluated system using AceForce 1.0. The green and yellow shaded areas represent absolute error thresholds of 1 kcal/mol and 2 kcal/mol, respectively. Additional metrics, including mean absolute error (MAE), root mean square error (RMSE), Kendall tau correlation ($\tau$), and top 30% and top 5 compound identification accuracy, are also displayed. 95% confidence interval values for the relevant metrics are shown in brackets.
3.2 Identifying the top compounds in a series
In benchmark studies, the primary objective is often to achieve the highest accuracy and minimize prediction errors. However, in practical drug discovery settings, such as hit-to-lead or lead optimization campaigns, the primary goal is to identify the most active compounds from a pool of candidates, even if their predicted affinities are somewhat overestimated. This prioritization aids in streamlining the selection process for synthesis or purchasing decisions.
To evaluate the practical utility in such scenarios, we assessed its ability to rank ligands correctly across different levels of priority. Specifically, we examined the top 5 ligands and the top 30% of ligands. The top 5 was measured by the Overlap Coefficient, which quantifies the proportion of correctly ranked ligands among the true top 5. The top 30% was assessed using the Fraction of Best Ligands (FBL) metric, which evaluates the proportion of ligands in the predicted top 30% that align with the true top 30%. This dual assessment allows us to determine not only the accuracy of identifying the very top-ranked ligands but also how well the method performs across a broader range of predicted positions.
Results in (Figure 5) indicate that AceForce 1.0 generally performs as well as or better than the other methods in identifying the most potent compounds. For example, in datasets such as BACE, CDK2, P38, and TYK2, AceForce 1.0 successfully identified at least 50% of the top 30% compounds (59%, 52%, 50%, and 84%, respectively), outperforming GAFF2 runs in all cases. However, the identification of top compounds by AceForce 1.0 is less effective in some datasets, such as JNK1, MCL1, and THROMBIN, where the top 30% identification is considerably lower. We believe that this metric is particularly valuable for larger datasets such as BACE, MCL1, and P38. For these datasets, AceForce 1.0 generally shows improved performance compared to GAFF2, though with variability across targets: BACE and P38 exhibit clear improvements, while MCL1 remains a challenge.
In terms of top 5 identification, the differences between methods were generally small. For MCL1, AceForce 1.0 placed one ligand in the top 5, while GAFF2 and OPLS4 identified two and three, respectively. In most other cases, the same number of top-ranked molecules were identified, with discrepancies of no more than a single ligand when comparing AceForce 1.0 to OPLS4 (Figure 5). Notably, even in cases where AceForce 1.0 exhibited higher overall error rates compared to OPLS4, its compound prioritization remained comparable.
Figure 5: Identification of top compounds across datasets. The left plot shows the accuracy in identifying compounds within the top 30%, while the right plot focuses on the top 5 compounds in each series. Results are compared across the evaluated methods: AceForce 1.0, GAFF2, and OPLS4 with FEP+.
3.3 Timesteps: 1fs vs 2fs
One of the key challenges of NNP/MM RBFE calculations is the increased computational cost required to execute these simulations. This is primarily due to two factors: the overhead associated with running the NNP and the limitations on the simulation timestep. While MM calculations can be confidently performed with a 4 fs timestep, NNP/MM runs have traditionally been restricted to 1 fs. Previous attempts to increase the timestep frequently resulted in simulation failures.
With the AceForce 1.0 model for TorchMD-Net, we observed improved stability at a 2fs timestep, significantly enhancing the computational speed of NNP/MM simulations. To evaluate whether this larger timestep maintains accuracy, we conducted RBFE calculations on a subset of the JACS dataset at a timestep of both 1 and 2fs. Our findings (Figure 6 and Supporting TablesS10 and S11) indicate that 2 fs timestep simulations demonstrate comparable accuracy to 1 fs runs for most targets. In these cases, RMSE and Kendall tau correlation values remained nearly identical. The impact on top compound identification is shown in Supporting Table S12. Considering the variability between runs, we can conclude that it is possible to run at simulations at 2 fs.
Figure 6: Comparison of AceForce 1.0 model accuracy at 1fs and 2fs timesteps. The top panel shows Root Mean Squared Error (RMSE) and the bottom panel shows Kendall tau correlation for $\Delta G$ values across a subset of systems. While runs at both timesteps yield similar performance, the 1fs timestep calculations maintain slightly higher accuracy.
3.4 Comparison with ANI-2x
We compare now the performance of AceForce to ANI-2x in terms of accuracy.22 From the JACS dataset, we computed several edges for CDK2, JNK1, P38, and TYK2. Due to limitations in ANI-2x’s atom type and charge handling, the datasets for JNK1 and CDK2 are not complete, so we compare only the same edges that both methods could calculate. In contrast, the P38 and TYK2 datasets are complete for both models, allowing for a more direct assessment of their performance. For this reason, a comparison between $\Delta G$ values is only done for P38 and TYK2. RMSE and MAE values for the computed $\Delta\Delta G$ edges are presented in Figure 7. Additionally, table S9 presents a summary of the $\Delta G$ results, while Table S8 provides the $\Delta\Delta G$ statistics for all four targets. For $\Delta\Delta G$ calculations on P38, AceForce achieves a slightly lower RMSE of 1.04 kcal/mol compared to ANI-2x’s 1.17 kcal/mol, alongside a lower MAE (0.84 vs. 0.91 kcal/mol). Additionally, AceForce shows a higher Kendall tau correlation (0.69 vs. 0.59). On TYK2, both models deliver similar accuracy, with AceForce showing a slightly lower RMSE (0.55 vs. 0.56 kcal/mol) and a modestly higher Kendall tau (0.73 vs. 0.67). For CDK2, AceForce yields an RMSE of 1.03 kcal/mol, slightly higher than ANI-2x’s 0.83 kcal/mol, and a Kendall tau of 0.46 compared to ANI-2x’s 0.62, indicating that ANI-2x captures the ranking of these particular ligands more effectively. Conversely, on JNK1, AceForce attains an RMSE of 0.92 kcal/mol vs. ANI-2x’s 0.90 kcal/mol, and a higher Kendall tau (0.46 vs. 0.43), suggesting a marginal ranking advantage for AceForce on this target. For absolute binding free energies ($\Delta G$), AceForce again demonstrates improved overall performance. On P38, AceForce achieves an RMSE of 0.77 kcal/mol and a Kendall tau of 0.72, and ANI-2x (RMSE 0.93, Kendall tau 0.58). Similarly, on TYK2, AceForce reaches an RMSE of 0.47 kcal/mol and Kendall tau of 0.85, slightly surpassing ANI-2x’s RMSE of 0.50 and Kendall tau of 0.82. Overall, these findings indicate that AceForce generally perform better across the evaluated systems with the additional benefit that it is not limited to those.
Figure 7: Comparison of AceForce and ANI-2x calculations for $\Delta\Delta G$ values.
3.5 Going beyond the limits of the model
PTP1B is a protein target included in the JACS dataset, characterized by ligands uniformly carrying a -2 charge. AceForce 1.0, however, has been trained on molecules with charges limited to -1, 0, and +1, making its applicability for this target troublesome. This limitation of AceForce will be solved in version 1.1. We conducted RBFE calculations for PTP1B to assess the model’s performance under such constraints. As anticipated, the predictions exhibited significant errors, with RMSE exceeding 3 kcal/mol for both $\Delta\Delta G$ and $\Delta G$ values and two out of the 49 runs crashing. Furthermore, the correlation performance was notably poor, reflected by a Kendall tau coefficient of -0.31, and the method failed to correctly identify any of the top ligands in the dataset (Figure8). This outcome aligns with expectations, emphasizing the need for further training of the model on an extended range of ligand charges to improve its generalization and predictive capabilities.
Figure 8: Top: Example on one of the ligands of the PTP1B dataset, highlighting the -1 charge at each of the two COO- groups. Bottom: scatterplot for the relative and absolute binding free energy calculations. Relevant metrics are reported on top of each plot
4 Conclusions
We are presenting our NNP/MM approach to perform RBFE calculations. We are using AceForce 1.0, a novel NNP model that supports a broader range of atom elements and charged molecules. To assess its accuracy, we benchmarked AceForce 1.0 using the JACS dataset27 (without PTP1B due to charge limitations) and compared it to an MM-based approach as well as the industry state-of-the-art alternative in OPLS4 with FEP+.28 AceForce 1.0 demonstrated overall improved performance to MM-based calculations like GAFF2. These improvements highlight NNP potentials as an accurate alternative to MM approaches for RBFE calculations. In comparison to OPLS4, the state-of-the-art approach in the industry, AceForce 1.0 provided competitive or superior correlation performance on half of the evaluated systems but underperformed on the rest. These results indicate that further training of the potential may be needed to consistently achieve SOTA accuracy across all targets. Additionally, when compared to ANI-2x, the initial NNP tested in our previous work,22 AceForce 1.0 delivered better accuracy overall.
One of the main limitations for NNPs remains computational speed, primarily due to timestep restrictions. To enhance computational efficiency, we demonstrated the feasibility of increasing the timestep from 1fs to 2 fs. This adjustment allows for stable simulations with similar accuracy in most cases. The results suggest that a 2 fs timestep can be used for more efficient drug discovery workflows.
Future work will focus on testing and developing new versions of AceForce with an expanded training dataset to enhance predictive accuracy and general applicability. We also aim to enable stable simulations at larger timesteps, potentially 3fs or even 4fs, to reduce the computational cost of NNP/MM calculations to levels comparable with MM-based methods. Additionally, we plan to expand RBFE studies to include other benchmark datasets.
5 Acknowledgement
The authors thank the volunteers of GPUGRID.net for donating computing time. This project has received funding from the Torres-Quevedo Programme from the Spanish National Agency for Research (PTQ2020-011145 / AEI / 10.13039/501100011033).
6 Conflict of interests
All the authors have a potential conflict of interest due to direct interests in Acellera Therapeutics (See affiliations).
7 Code and Model availability
AceForce models are available from HuggingFace https://huggingface.co/Acellera/AceForce .
Examples of how to use it are available: 1) Run ML potential molecular simulations of a small molecule using ACEMD https://software.acellera.com/acemd/nnp.html. 2) For a tutorial on running mixed protein-ligand simulations, refer to NNP/MM https://software.acellera.com/acemd/nnpmm.html.
A tutorial on how to run the RBFE calculations, input data and structures are available at https://github.com/Acellera/quantumbind_rbfe.